# CounselBench: A Large-Scale Expert Evaluation and Adversarial Benchmarking of Large Language Models in Mental Health Question Answering

### v2

@@ Abstract
Medical question answering (QA) benchmarks often focus on multiple-choice or fact-based tasks, leaving open-ended answers to real patient questions underexplored. This gap is particularly critical in mental health, where patient questions often mix symptoms, treatment concerns, and emotional needs, requiring answers that balance clinical caution with contextual sensitivity. We present CounselBench, a large-scale benchmark developed with 100 mental health professionals to evaluate and stress-test large language models (LLMs) in realistic help-seeking scenarios. The first component, CounselBench-EVAL, contains 2,000 expert evaluations of answers from GPT-4, LLaMA 3, Gemini, and human therapists on patient questions from the public forum CounselChat. Each answer is rated across six clinically grounded dimensions, with span-level annotations and written rationales. Expert evaluations show that while LLMs achieve high scores on several dimensions, they also exhibit recurring issues, including unconstructive feedback, overgeneralization, and limited personalization or relevance. Responses were frequently flagged for safety risks, most notably unauthorized medical advice. Follow-up experiments show that LLM judges systematically overrate model responses and overlook safety concerns identified by human experts. To probe failure modes more directly, we construct CounselBench-Adv, an adversarial dataset of 120 expert-authored mental health questions designed to trigger specific model issues. Evaluation of 3,240 responses from nine LLMs reveals consistent, model-specific failure patterns. Together, CounselBench establishes a clinically grounded framework for benchmarking LLMs in mental health QA.
@@ 1 Introduction
These challenges are particularly critical in mental health QA. Here, patient queries are expressed as free-text descriptions of symptoms, treatment concerns, and emotional needs, and helpful responses must balance empathy, actionable guidance, and professional boundaries (Elliott et al., 2011; Na, 2024; Meadi et al., 2025). Unlike other medical domains, where diagnostic tests or structured knowledge often constrain the answer space, mental health QAs are inherently subjective and context-dependent (Association, 2013; Hormazábal-Salgado et al., 2024). Moreover, mental health QA has direct real-world significance: platforms such as CounselChat (Bertagnolli, 2020), peer-support forums (Pendse et al., 2019), EHR-based messaging, and digital mental health services like NOCD (NOCD, 2018) provide brief, asynchronous answers at scale. For many individuals, these platforms are a primary source of support (Naslund et al., 2016; Merchant et al., 2022). In such settings, even small errors, such as speculative claims, unsafe recommendations, or dismissive tone, can have immediate negative consequences. Evaluating LLMs on authentic, open-ended mental health QA is therefore both technically and socially urgent.
@@
    CounselBench-Eval: we recruit 100 licensed mental health professionals to provide 2,000 evaluations of answers to real patient questions, with span-level annotations and written rationales, making it among the largest expert-annotated datasets for mental health QA.
@@
    CounselBench-Adv: we construct an adversarial dataset of 120 expert-authored mental health questions designed to elicit systematic failure modes, providing targeted signals for model improvement.
@@ 2 Related Work
LLMs in Medical QA Benchmarks. Early medical QA benchmarks emphasized closed-form accuracy, such as multiple-choice exams (Pal et al., 2022) or retrieval-based tasks (Zhu et al., 2020). Recent benchmarks broaden evaluation with multi-axis rubrics. For example, MultiMedQA (Singhal et al., 2022) introduces physician-designed criteria such as factuality and bias. HealthBench (Arora et al., 2025) further scales to tens of thousands of physician-curated items in general medical domains. These works advance evaluation beyond accuracy but remain focused on structured medical knowledge. In contrast, our work targets authentic mental health QA, where
Evaluating Open-Ended Medical QA. Evaluating free-form medical QA remains difficult. Automated metrics (e.g., BLEU) scale easily but poorly capture qualities like specificity or safety (Croxford et al., 2024). Some studies incorporate expert review, but often with small panels due to cost (Hosseini et al., 2024; Kim et al., 2024; Manes et al., 2024). LLM-as-judge approaches are emerging (Szymanski et al., 2024), yet their reliability in clinical domains requires human validation. Many works also use non-experts or heuristics (Turcan & McKeown, 2019), risking inconsistent results.
Benchmarks for Mental Health QA. Existing mental health QA datasets mostly focus on exam-style knowledge tests with objective answer keys (Racha et al., 2025; Nguyen et al., 2025). Others adopt psychological rubrics (Li et al., 2024; Chiu et al., 2024) but rarely involve clinicians in design, often relying on crowdworkers or heuristics (Young et al., 2024). Our evaluation protocol is co-designed with clinical psychologists and scaled annotations to 100 licensed or trained professionals, ensuring clinical validity and reproducibility.
Adversarial Evaluation in Mental Health QA. Prior studies have highlighted risks of LLMs in mental health support, such as unsafe advice or discrepant responses based on inferred demographics (Lawrence et al., 2024; Imran et al., 2023; Gabriel et al., 2024). However, most findings come from post hoc analyses of deployed systems (Elyoseph & Levkovich, 2023), which reveal problems retrospectively rather than prospectively stress-testing models. A few recent works adapt red-teaming to mental health (Grabb et al., 2024; Schoene & Canca, 2025), but their failure modes are predefined and literature-driven, limiting coverage of the diverse failure patterns that emerge in practice. In contrast, our adversarial dataset is empirically grounded: 10 licensed mental health professionals authored 120 prompts from failures observed in CounselBench-Eval, creating an expert-crafted benchmark for prospectively eliciting clinically meaningful vulnerabilities.
@@ 3 Overview of CounselBench
  • COUNSELBENCH-Eval: a benchmark of 2,000 expert evaluations of LLMs and human responses to real patient queries, with numeric scores, span-level annotations, and rationales.
  • COUNSELBENCH-ADV: an adversarial dataset of 120 expert-authored questions designed to expose identified failure modes in LLMs.
Figure 1: Overview of CounselBench benchmark. COUNSELBENCH-EVAL (left) includes expert evaluation of LLMs and human responses to real counseling questions. COUNSELBENCH-ADV (right) includes adversarial questions authored by clinicians to target identified LLM failure modes. See Appendix B for license/degree types and specialization areas.
Research Pipeline. We began by sampling 100 real patient questions from the public forum CounselChat (Bertagnolli, 2020), spanning 20 common mental health topics (§4.1). Each question was answered by GPT-4, Gemini, LLaMA-3, and a human therapist (§4.2). To evaluate responses, we defined six evidence-based dimensions grounded in literature: overall quality, empathy, specificity, factual consistency, medical advice, and toxicity (§4.3). We then recruited 100 mental health professionals for evaluation (§4.4). Besides assigning ratings, annotators provided span-level annotations and written rationales, yielding rich insights into response quality and safety. These evaluations form the COUNSELBENCH-EVAL dataset. We also prompted nine LLMs to rate the same responses under the same rubric (§4.6), enabling systematic comparison between human and LLM judges. To probe safety directly, 10 experts authored 120 adversarial questions, creating the CounselBench-Adv dataset (§5), to trigger failure patterns identified in CounselBench-Eval.
@@ 4 CounselBench-EVAL: Expert Evaluation of LLMs for Open‑Ended Mental‑Health QA
We develop COUNSELBENCH-EVAL to systematically assess LLM behavior on authentic mental health questions. This section describes the construction pipeline: real-world mental health QA sourcing (§4.1), response generation (§4.2), rubric design (§4.3), human annotation protocol (§4.4), results from expert scoring and analysis (§4.5), and experiments with LLM-based evaluators (§4.6).
@@ 4.1 Real-World Mental Health Question and Answering
We source mental health QA pairs from CounselChat, an online platform where licensed therapists publicly respond to anonymous patient questions (Bertagnolli, 2020). All posts were authored before 2022, ensuring that all human responses predate the widespread use of LLMs like ChatGPT.
Question Curation and Topic Coverage. To ensure both topical diversity and response quality, we selected the top 20 topics based on predefined criteria (Appendix  LABEL:appendix_question_filtering). From each topic, we sampled five questions, resulting in 100 questions in total. The topics were: (1) depression, (2) relationships, (3) anxiety, (4) family conflict, (5) parenting, (6) self esteem, (7) relationship dissolution, (8) behavioral change, (9) anger management, (10) trauma, (11) marriage, (12) domestic violence, (13) grief and loss, (14) social relationships, (15) workplace relationships, (16) legal regulatory, (17) substance abuse, (18) counseling fundamentals, (19) eating disorders, (20) professional ethics.
Human Answers. Although all answers on CounselChat were written by certified clinicians, forum contributions are informal and vary in quality. To reduce this variability and focus on consistently well-regarded content, we selected the top-voted answer for each question (Appendix  LABEL:appendix_human_responses). Upvotes are a standard, publicly observable indicator of perceived helpfulness and relevance in peer-support forums, and have been widely used in prior work to capture perceived helpfulness(Jiang et al., 2020; Huang et al., 2025), social support (De Choudhury & De, 2014), engagement (Gao et al., 2020), community reception (Alghamdi et al., 2025), and narrative quality (Wang et al., 2017).
@@ 4.2 Generating Counseling Responses from LLMs
Pilot Study. We first piloted domain-specific models such as MentalLLaMA (Yang et al., 2023) and Meditron (Chen et al., 2023b). As shown in Appendix LABEL:sec:domain-specific-experiments, these models performed poorly on open-ended mental health QA, likely due to narrow fine-tuning or lack of instruction tuning. We therefore focused on general-purpose LLMs that represent current deployment paradigms
Model Selection. To ensure fair comparison and reduce annotator fatigue, we limited our evaluation to three LLMs per question. Each mental health professional blindly annotated four answers side by side (three LLM-generated and one human-written), enabling consistent evaluation while keeping the task cognitively manageable. The selected models are: (1) GPT-4-0613 (via OpenAI API), a widely adopted commercial model, (2) LLaMA-3.3-70B-Instruct (via Hugging Face), a leading open-source, instruction-tuned model from Meta, (3) Gemini-1.5-Pro (via Google Generative AI SDK), which powers responses in Google Search and Assistant. Together, they represent major deployment paradigms: commercial APIs, open-source checkpoints, and search-integrated platforms. Each model is prompted with the original patient question and asked to respond. Full prompt templates and generation details are provided in Appendix D.1.
@@ 4.3 Evaluation Rubric and Paradigms
Rating Rationale & Text Extraction To capture the context and reasoning behind human evaluations of counseling messages, we asked annotators (§4.4) to provide open-text explanations for their ratings on Overall Quality, Medical Advice, Factual Consistency, and Toxicity. Additionally, we asked them to identify and extract specific portions of each response that supported their ratings. Two researchers on the team checked and verified the quality of text responses after data collection. In general, human annotators provided detailed written feedback, with an median length of 576.5 written words (IQR: 339.8 - 866.5) per survey combining all rationales.
@@ 4.4 Expert Annotation Protocol
Annotation Protocol. Each annotator was assigned a randomly sampled survey containing 5 counseling questions, each paired with 4 responses: one from an online human therapist and three from different LLMs. The four responses for each question are in randomized order to mitigate position bias. This resulted in 20 QA pairs per annotator. Each pair was independently evaluated by five different professionals, supporting inter-rater reliability and reducing individual bias. For each question, all four responses were rated by the same group of annotators, enabling direct and fair comparison across models. In total, the study includes 2,000 annotated QA pairs. Annotators were blind to the source of each response and were not informed that some had been generated by LLMs.
Annotator Demographics and Engagement Collectively, the 100 annotators held 32 distinct types of licenses and degrees, and reported expertise spanning 43 specialized counseling areas (summarized in Figure 2; more details in Appendix B). This broad coverage ensures that our evaluation reflects a diverse and representative cross-section of real-world practice. The demographic composition of our annotator sample aligns with national statistics for the U.S. counseling profession, which shows a higher proportion of white and female practitioners (McGuire & Miranda, 2008; Verdieu, 2024; American Psychological Association, 2022). The median written rationale length was 576.5 words across 20 QA pairs (IQR: 339.8 - 866.5), and the median time spent was 1 hr 22 min (IQR: 52 min - 3 hrs 9 min), indicating substantial engagement.
4.5 Expert Evaluation Results: Counseling Quality Across LLMs and Humans
@@
Annotation Analysis. Additional analyses of annotator behavior and score distributions are reported in Appendix C. We find that annotators who spent more time (top 50%) wrote longer rationales and labeled more sentences as Medical Advice. Significant differences in Empathy, Medical Advice, and Factual Consistency were also observed across annotators with different years of experience. Finally, Table LABEL:tab:uncertain-percentage reports the fraction of “I am not sure” selections, which were infrequent overall.
Qualitative Analysis of Annotator Rationales To better understand expert judgments, we conducted a thematic analysis following established protocol (Clarke & Braun, 2017). Two researchers on our team independently reviewed and thematically coded annotators’ open-text explanations. Seven primary issues among low-quality responses are identified (listed in Appendix LABEL:sub_appendeix_llm_judge_eval). We then prompted GPT-4.1 to label all low-rated responses (overall score $\leq$ 2) using these categories. Each response could exhibit multiple issues. A detailed breakdown is provided in Figure LABEL:fig:low_overall (Appendix LABEL:sub_appendeix_llm_judge_eval). The top two most frequently flagged failure types for each model are:
@@
Clinical Boundaries: Unauthorized Medical Advice in LLM responses To further examine safety risks, we analyzed the response excerpts flagged by expert annotators for containing medical advice. The most concerning identified issue was the inclusion of unauthorized clinical recommendations, especially prescribed psychotropic treatments (i.e., medications for mental health treatments, according to National Institute of Mental Health (of Mental Health, 2025)).

### v3

@@ Abstract
Medical question answering (QA) benchmarks often focus on multiple-choice or fact-based tasks, leaving open-ended answers to real patient questions underexplored. This gap is particularly critical in mental health, where patient questions often mix symptoms, treatment concerns, and emotional needs, requiring answers that balance clinical caution with contextual sensitivity. We present CounselBench, a large-scale benchmark developed with 100 mental health professionals to evaluate and stress-test large language models (LLMs) in realistic help-seeking scenarios. The first component, CounselBench-EVAL, contains 2,000 expert evaluations of answers from GPT-4, LLaMA 3, Gemini, and online human therapists on patient questions from the public forum CounselChat. Each answer is rated across six clinically grounded dimensions, with span-level annotations and written rationales. Expert evaluations show that while LLMs achieve high scores on several dimensions, they also exhibit recurring issues, including unconstructive feedback, overgeneralization, and limited personalization or relevance. Responses were frequently flagged for safety risks, most notably unauthorized medical advice. Follow-up experiments show that LLM judges systematically overrate model responses and overlook safety concerns identified by human experts. To probe failure modes more directly, we construct CounselBench-Adv, an adversarial dataset of 120 expert-authored mental health questions designed to trigger specific model issues. Expert evaluation of 1,080 responses from nine LLMs reveals consistent, model-specific failure patterns. Together, CounselBench establishes a clinically grounded framework for benchmarking LLMs in mental health QA.
@@ 1 Introduction
These challenges are particularly critical in mental health QA. Here, patient queries are expressed as free-text descriptions of symptoms, treatment concerns, and emotional needs, and helpful responses must balance empathy, actionable guidance, and professional boundaries (Elliott et al., 2011; Na, 2024; Meadi et al., 2025). Unlike other medical domains, where diagnostic tests or structured knowledge often constrain the answer space, mental health QAs are inherently subjective and context-dependent (Association, 2013; Hormazábal-Salgado et al., 2024). Moreover, mental health QA has direct real-world significance: platforms such as CounselChat (Bertagnolli, 2020), peer-support forums (Pendse et al., 2019), EHR-based messaging, and digital mental health services like NOCD (NOCD, 2018) primarily operate through single-turn interactions, with limited follow-up in most cases. These setting are designed for brief, asynchronous exchanges that offer emotional support, psychoeducation, or coping strategies in a concise format. Such platforms are widely used and often serve as a primary source of support, especially for individuals who may not have access to or feel ready for full-length therapy (Naslund et al., 2016; Merchant et al., 2022). In such settings, even small errors, such as speculative claims, unsafe recommendations, or dismissive tone, can have immediate negative consequences. Evaluating LLMs on authentic, open-ended mental health QA in a standardized framework that enables consistent comparison of model responses to identical patient questions is therefore both technically and socially urgent.
@@
    CounselBench-Eval: we recruit 100 licensed or professionally trained mental health professionals to provide 2,000 evaluations of answers to real patient questions. These evaluations cover 4 answer providers (GPT-4, LLaMA-3, Gemini, and online human therapists) across 100 questions, with 5 independent expert annotations per response. Each annotation includes span-level labels and written rationales.
@@
    CounselBench-Adv: we construct an adversarial dataset of 120 expert-authored mental health questions designed to elicit systematic failure modes. For these questions, we collect 1,080 responses from nine models and have them evaluated by professionals, which identifies targeted signals for model improvement.
@@ 2 Related Work
LLMs in Medical QA Benchmarks. Early medical QA benchmarks emphasized closed-form accuracy, such as multiple-choice exams (Pal et al., 2022) or retrieval-based tasks (Zhu et al., 2020). Recent benchmarks broaden evaluation beyond accuracy with multi-axis rubrics. For example, MultiMedQA (Singhal et al., 2022) introduces physician-designed criteria such as factuality and bias. HealthBench (Arora et al., 2025) further scales to tens of thousands of physician-curated items in general medical domains. However, they remain focused on structured medical knowledge or rely on complex schemes that are difficult to apply in practice. In contrast, our work targets authentic mental health QA, where patient-facing questions require safety, empathy, and contextual sensitivity.
Evaluating Open-Ended Medical QA. Evaluating free-form medical QA remains difficult. Automated metrics (e.g., BLEU) scale easily but poorly capture qualities like specificity or safety (Croxford et al., 2024). Some studies incorporate expert review, but often with small panels due to cost (Hosseini et al., 2024; Kim et al., 2024; Manes et al., 2024). LLM-as-judge approaches are emerging (Szymanski et al., 2024), yet their reliability in healthcare domains requires human validation.
Benchmarks for Mental Health QA. Existing mental health QA datasets mostly focus on exam-style knowledge tests with objective answer keys (Racha et al., 2025; Nguyen et al., 2025). Others adopt psychological rubrics (Li et al., 2024), but rarely involve clinicians in design (Chiu et al., 2024); the evaluation is often scaled by non-expert raters or heuristics (Young et al., 2024). Our evaluation protocol is co-designed with clinical psychologists and scaled annotations to 100 licensed or trained professionals, ensuring clinical validity and reproducibility.
Adversarial Evaluation in Mental Health QA. Prior studies have highlighted risks of LLMs in mental health support, such as unsafe advice or discrepant responses based on inferred demographics (Lawrence et al., 2024; Imran et al., 2023; Gabriel et al., 2024). However, most findings come from post hoc analyses of deployed systems (Song et al., 2025; Li et al., 2025b), which reveal problems retrospectively rather than prospectively stress-testing models. A few recent works adapt red-teaming to mental health (Grabb et al., 2024; Schoene & Canca, 2025), but their failure modes are predefined and literature-driven, limiting coverage of the diverse failure patterns that emerge in practice. In contrast, our adversarial dataset is empirically grounded: 10 mental health professionals authored 120 prompts from failures observed in CounselBench-Eval, creating an expert-crafted benchmark for prospectively eliciting clinically meaningful vulnerabilities.
@@ 3 Overview of CounselBench
  • COUNSELBENCH-Eval: a benchmark of 2,000 expert evaluations of LLMs and online human therapist responses to real patient queries, with numeric scores, span-level annotations, and rationales.
  • COUNSELBENCH-ADV: an adversarial dataset of 120 expert-authored questions designed to expose identified failure modes in LLMs, paired with 1,080 responses from 9 LLMs, each annotated for whether the targeted failure mode is present.
Figure 1: Overview of CounselBench benchmark. COUNSELBENCH-EVAL (left) includes expert evaluation of LLMs and online human therapist responses to real counseling questions. COUNSELBENCH-ADV (right) includes adversarial questions authored by clinicians to target identified LLM failure modes. See Appendix B for license/degree types and specialization areas.
Research Pipeline. We began by selecting 100 real patient questions from the public forum CounselChat (Bertagnolli, 2020), spanning 20 common mental health topics (§4.1). Each question was answered by GPT-4, Gemini, LLaMA-3, and an online human therapist (§4.2). To evaluate responses, we defined six evidence-based dimensions grounded in literature: overall quality, empathy, specificity, factual consistency, medical advice, and toxicity (§4.3). We then recruited 100 mental health professionals for evaluation (§4.4). Besides assigning ratings, annotators provided span-level annotations and written rationales, yielding rich insights into response quality and safety. These evaluations form the COUNSELBENCH-EVAL dataset. We also prompted nine LLMs to rate the same responses under the same rubric (§4.6), enabling systematic comparison between human and LLM judges. To probe safety directly, 10 experts authored 120 adversarial questions and additional 5 expert annotated their identified failure modes, creating the CounselBench-Adv dataset (§5), to trigger failure patterns identified in CounselBench-Eval.
@@ 4 CounselBench-EVAL: Expert Evaluation of LLMs for Open‑Ended Mental‑Health QA
We develop COUNSELBENCH-EVAL to systematically assess LLM behavior on authentic mental health questions. This section describes the construction pipeline: real-world mental health QA sourcing (§4.1), response generation (§4.2), rubric design (§4.3), human annotation protocol (§4.4), results from expert scoring and analysis (§4.5), and experiments with LLM judges (§4.6).
@@ 4.1 Real-World Mental Health Question and Answering
We source mental health QA pairs from CounselChat, a platform where licensed therapists publicly respond to anonymous patient questions (Bertagnolli, 2020). All posts were authored before 2022, ensuring that all online therapist responses predate the widespread use of LLMs like ChatGPT.
Question Curation and Topic Coverage. To ensure both topical diversity and response quality, we selected the topics and questions based on predefined criteria (Appendix  LABEL:appendix_question_filtering). From each topic, we selected five questions with the highest response upvotes, resulting in 100 questions in total. The topics were: (1) depression, (2) relationships, (3) anxiety, (4) family conflict, (5) parenting, (6) self esteem, (7) relationship dissolution, (8) behavioral change, (9) anger management, (10) trauma, (11) marriage, (12) domestic violence, (13) grief and loss, (14) social relationships, (15) workplace relationships, (16) legal regulatory, (17) substance abuse, (18) counseling fundamentals, (19) eating disorders, (20) professional ethics. The keywords for each question are summarized in Table LABEL:tab:keyword_per_question.
Human Answers. All responses on CounselChat are written by verified, non-anonymous therapists with public profiles; only patient questions are anonymous (Bertagnolli, 2020). Despite that, forum contributions are informal and vary in quality. To reduce this variability and focus on consistently well-regarded content, we selected the top-voted answer for each question (Appendix  LABEL:appendix_human_responses). Upvotes are a standard, publicly observable indicator of perceived helpfulness and relevance in peer-support forums, and have been widely used in prior work to capture perceived helpfulness(Jiang et al., 2020; Huang et al., 2025), social support (De Choudhury & De, 2014), engagement (Gao et al., 2020), community reception (Alghamdi et al., 2025), and narrative quality (Wang et al., 2017).
@@ 4.2 Generating Counseling Responses from LLMs
Pilot Study. We first piloted domain-specific models such as MentalLLaMA (Yang et al., 2023) and Meditron (Chen et al., 2023b). As shown in Appendix LABEL:sec:domain-specific-experiments, these models performed poorly on open-ended mental health QA, likely due to narrow fine-tuning or lack of instruction tuning. We therefore focused on general-purpose LLMs that represent current deployment paradigms .
Model Selection. To ensure fair comparison and reduce annotator fatigue, we limited our evaluation to three LLMs per question. Each mental health professional blindly annotated four answers (three LLM-generated and one human-written), enabling consistent evaluation while keeping the task cognitively manageable. The selected models are: (1) GPT-4-0613 (via OpenAI API), a widely adopted commercial model, (2) LLaMA-3.3-70B-Instruct (via Hugging Face), a leading open-source, instruction-tuned model from Meta, (3) Gemini-1.5-Pro (via Google Generative AI SDK), which powers responses in Google Search and Assistant. Together, they represent major deployment paradigms: commercial APIs, open-source checkpoints, and search-integrated platforms. Each model is prompted with the original patient question and asked to respond. Full prompt templates and generation details are provided in Appendix D.1.
@@ 4.3 Evaluation Rubric and Paradigms
Rating Rationale & Text Extraction To capture the context and reasoning behind human evaluations of mental health QA, we asked annotators (§4.4) to provide open-text explanations for their ratings on Overall Quality, Medical Advice, Factual Consistency, and Toxicity. Additionally, we asked them to identify and extract specific portions of each response that supported their ratings. Two researchers on the team checked and verified the quality of text responses after data collection. In general, human annotators provided detailed written feedback, with an median length of 576.5 written words (IQR: 339.8 - 866.5) per survey combining all rationales.
@@ 4.4 Expert Annotation Protocol
Annotation Protocol. Each annotator was assigned a randomly sampled survey containing 5 counseling questions, each paired with 4 responses: one from an online human therapist and three from different LLMs. The four responses for each question are in randomized order to mitigate position bias. This resulted in 20 QA pairs per annotator. Each pair was independently evaluated by five different professionals, supporting inter-rater reliability and reducing individual bias. For each question, all four responses were rated by the same group of annotators, enabling direct and fair comparison across models. In total, the study includes 2,000 annotated QA pairs (100 questions × 4 answers × 5 annotations). Annotators were blind to the source of each response and were not informed that some had been generated by LLMs.
Annotator Demographics and Engagement. Collectively, the 100 annotators held 32 distinct types of licenses and degrees, and reported expertise spanning 43 specialized counseling areas (summarized in Figure 2; more details in Appendix B). This broad coverage ensures that our evaluation reflects a diverse and representative cross-section of real-world practice. The demographic composition of our annotator sample aligns with national statistics for the U.S. counseling profession, which shows a higher proportion of white and female practitioners (McGuire & Miranda, 2008; Verdieu, 2024; American Psychological Association, 2022). The median written rationale length was 576.5 words across 20 QA pairs (IQR: 339.8 - 866.5), and the median time spent was 1 hr 22 min (IQR: 52 min - 3 hrs 9 min), indicating substantial engagement.
4.5 Expert Evaluation Results: Counseling Quality Across LLMs and Online Human Therapists
@@
Annotation Analysis. Additional analyses of annotator behavior and score distributions are reported in Appendix C. We find that annotators who spent more time (top 50%) wrote longer rationales and those in the top 25% of time spent also labeled more sentences as Medical Advice. Significant differences in Empathy, Medical Advice, and Factual Consistency were also observed across annotators with different years of experience. Finally, Table LABEL:tab:uncertain-percentage reports the fraction of “I am not sure” selections, which were infrequent overall.
Qualitative Analysis of Annotator Rationales. To better understand expert judgments, we conducted a thematic analysis following established protocol (Clarke & Braun, 2017). Two researchers on our team independently reviewed and thematically coded annotators’ open-text explanations. Seven primary issues among low-quality responses are identified (listed in Appendix LABEL:sub_appendeix_llm_judge_eval). We then prompted GPT-4.1 to label all low-rated responses (overall score $\leq$ 2) using these categories. Each response could exhibit multiple issues. A detailed breakdown is provided in Figure LABEL:fig:low_overall (Appendix LABEL:sub_appendeix_llm_judge_eval). The top two most frequently flagged failure types for each model are:
@@
Clinical Boundaries: Unauthorized Medical Advice in LLM responses. To further examine safety risks, we analyzed the response excerpts flagged by expert annotators for containing medical advice. The most concerning identified issue was the inclusion of unauthorized clinical recommendations, especially prescribed psychotropic treatments (i.e., medications for mental health treatments, according to National Institute of Mental Health (of Mental Health, 2025)).